VACUUM UV LIF MEASUREMENTS OF RO-VIBRATIONALLY EXCITED HYDROGEN MOLECULES IN A SUPERSONIC PLASMA EXPANSION

P. VANKAN, S. B. S. HEIL, D.C. SCHRAM, and \underline{R. ENGELN}, Department of Applied Physics, Centre for Plasma Physics and Radiation Technology (CPS), Eindhoven University of Technology, P.O. Box 513, NL-5600 MB Eindhoven, The Netherlands.

Rotationally and vibrationally excited hydrogen molecules are measured using Laser Induced Fluorescence (LIF) spectroscopy in the vacuum-UV. The laser radiation is generated via the Stimulated Anti-Stokes Raman Scattering (SARS) process, and is tunable between 120 nm and 230 nm. Up to 10 Anti-Stokes beams are simultaneously directed into the hydrogen plasma expansion. This so-called multiplex excitation makes an accurate wavelength calibration of the pump laser imperative. This calibration is performed by simultaneously recording the VUV-LIF signal from the plasma expansion and the well-known two-photon laser induced fluorescence spectrum of nitric oxide in a gas cell^{a}. The density of hydrogen molecules in the ro-vibrationally excited energy levels is calibrated on the basis of previously performed Coherent Anti-Stokes Raman Scattering measurements. Population distributions are measured in an expanding plasma to determine the production mechanism of the ro-vibrationally excited hydrogen molecules. In this contribution we present results on the population distribution of H_{2}^{r,v} from v = 2 to v = 6 in the electronic ground state. The ro-vibrational population distribution in the expanding hydrogen plasma can be described with a two temperature Boltzmann distribution. The population distribution in the low rotational levels (J ≤ 5) in the measured vibrational states can be described with a temperature of 800 K, while the higher rotational levels are populated according to a vibrational-like temperature of 3600K. In the v = 2 state rotational excitation is observed up to J = 19. A possible generation mechanism for the super-rotational excitation is the association of atomic hydrogen at the reactor wall.

aP. Vankan, S. B. S. Heil, S. Mazouffre, R. Engeln, D. C. Schram, H. F. Döbele, accepted for publication in Rev. Sci. Instrum., issue April 2004